Assume a human body has 100W(J/sec) of baseline metabolism rate (BMR), 1.5m^2 of surface area, and emissivity of 0.9. Note that for homeostasis the thermal energy of BMR must be balanced with an equal loss of radiation energy to the environment. Ignore the convection in this problem. If one layer of cloth that has an emissivity of 0.9 as well is covered on the skin (34C), How thick a layer will we need of for the thermal conductivity, k, equal to 0.25? The temperature of the outside surface of clothing is 31.3C.

The final answer should be 1.00cm thick, please show all your work, thank you!
Answer #1

Given,

P = 100 W ; A = 1.5 m^2 ; k = 0.25

Th = 34 deg C ; Tc = 0.25

We know that

P = Q/t = k A (Th - Tc)/d

d = k A (Th - Tc)/P

d = 0.25 x 1.5 (34 - 31.3)/100 = 0.01 m = 1 cm

Hence, d = 1 cm